# How Much is Enough? The Diminishing Returns of Tokenization Training Data

### v2

@@ Abstract
Tokenization, a crucial initial step in natural language processing, is governed by several key parameters, such as the tokenization algorithm, vocabulary size, pre-tokenization strategy, inference strategy, and training data corpus. This paper investigates the impact of an often-overlooked hyperparameter, tokenizer training data size. We train BPE, UnigramLM, and WordPiece tokenizers across various vocabulary sizes using English training data ranging from 1GB to 900GB. Our findings reveal diminishing returns as training data size increases beyond roughly 150GB, suggesting a practical limit to the improvements in tokenization quality achievable through additional data. We analyze this phenomenon and attribute the saturation effect to constraints introduced by the pre-tokenization stage. We then demonstrate the extent to which these findings can generalize by experimenting on data in Russian, a language typologically distant from English. While the limit appears to materialize at a later phase of pre-training, around 200GB, it is in fact observed. These results provide valuable insights for optimizing the tokenization process by reducing the compute required for training on large corpora and suggest promising directions for future research in tokenization algorithms.
@@ 2.2 Intrinsic analysis
Encourages optimal encoding by evaluating the ratio of the Shannon entropy to the maximum possible entropy of the token distribution. Higher entropy indicates a more diverse and unpredictable distribution, penalizing vocabularies which are skewed towards very frequent and/or very rare tokens.
@@ 3 The Limiting Role of Pre-Tokenization
4 Generalizing Over Languages
To better understand whether the impact of tokenizer training data on tokenizer characteristics discussed in this work is robust, we augment our analysis of English data with a focused case study on Russian. Unlike English, Russian is a morphologically rich language, characterized by extensive inflectional and derivational morphology. This makes it a compelling test case for evaluating how tokenizers handle linguistic complexity under conditions markedly different than those in English.
@@ 4.1 Vocabulary analysis
As depicted in Figure 8, the proportion of shared vocabulary between tokenizers and a reference tokenizer trained with 450GB of data consistently increases with larger Russian training datasets. When juxtaposed with English, we can infer a consistent trend across languages. However, a notable difference is in the rate of increase between the two languages. The rate at which the fraction of common vocabulary increases is comparatively lower for Russian. This could be attributed to the inherently more complex nature of the language’s morphology, where individual types of (mostly) nouns, verbs, and adjectives appear much less frequently than their English equivalents where one form subsumes several, or even dozens, inflected Russian forms.
@@ C.3 Computing Domain-Specific and General English Terms
We curate two text corpora: a domain corpus ($D$), which is a compilation of all our downstream domain-specific evaluation data, and a general corpus ($G$) [^15] https://huggingface.co/datasets/wikimedia/wikipedia, representing standard English usage. Both corpora are matched in size to ensure a fair comparison.

### v3

@@ Abstract
Tokenization, a crucial initial step in natural language processing, is governed by several key parameters, such as the tokenization algorithm, vocabulary size, pre-tokenization strategy, inference strategy, and training data corpus. This paper investigates the impact of an often-overlooked hyperparameter, tokenizer training data size. We train BPE, UnigramLM, and WordPiece tokenizers across various vocabulary sizes using English training data ranging from 1GB to 900GB. Our findings reveal diminishing returns as training data size increases beyond roughly 150GB, suggesting a practical limit to the improvements in tokenization quality achievable through additional data. We analyze this phenomenon and attribute the saturation effect to constraints introduced by the pre-tokenization stage. We then demonstrate the extent to which these findings can generalize by experimenting on data in Russian, a language typologically distant from English. For Russian text, we observe diminishing returns after training a tokenizer from 200GB of data, which is approximately 33% more than when training on English. These results provide valuable insights for optimizing the tokenization process by reducing the compute required for training on large corpora and suggest promising directions for future research in tokenization algorithms.
@@ 2.2 Intrinsic analysis
Encourages optimal encoding by evaluating the ratio of the Shannon entropy to the maximum possible entropy of the token distribution. Higher entropy indicates a more diverse and unpredictable distribution, penalizing vocabularies which are skewed towards very frequent and/or very rare tokens. We use $\alpha$ = 2.5 for our analysis following  Zouhar et al. (2023).
@@ 3 The Limiting Role of Pre-Tokenization
4 Generalizing To Russian
To better understand whether the impact of tokenizer training data on tokenizer characteristics discussed in this work is robust, we augment our analysis of English data with a focused case study on Russian. Compared to English, Russian is a more morphologically rich language, characterized by extensive inflectional and derivational morphology. This makes it a compelling test case for evaluating how tokenizers handle linguistic complexity under conditions markedly different than those in English.
@@ 4.1 Vocabulary analysis
As depicted in Figure 8, the proportion of shared vocabulary between tokenizers and a reference tokenizer trained with 450GB of data consistently increases with larger Russian training datasets. When juxtaposed with English, we can infer a consistent trend across languages. However, a notable difference is in the rate of increase between the two languages. The rate at which the fraction of common vocabulary increases is comparatively lower for Russian. This could be attributed to the comparatively more complex nature of the language’s morphology, where individual types of (mostly) nouns, verbs, and adjectives appear much less frequently than their English equivalents where one form subsumes several, or even dozens, inflected Russian forms.
@@ C.3 Computing Domain-Specific and General English Terms
We curate two text corpora: a domain corpus ($D$), which is a compilation of all our downstream domain-specific evaluation data, and a general corpus ($G$), [^15] https://huggingface.co/datasets/wikimedia/wikipedia representing standard English usage. Both corpora are matched in size to ensure a fair comparison.